OpenStudy (stamp):

[DiffEq] Find the general solution to the first order linear differential equation. \[(x^4+2y) dx - x dy = 0\]

OpenStudy (sshayer):

\[x^4+2y=x \frac{ dy }{ dx }\] \[\frac{ dy }{ dx }-\frac{ 2 }{ x }y=x^3\] \[I.F.=e ^{\int\limits \frac{ -2 }{ x }dx}=e ^{-2\ln x}=e ^{\ln x ^{-2}}=x ^{-2}=\frac{ 1 }{ x^2 }\] \[C.S.~is~y \times \frac{ 1 }{ x^2 }=\int\limits x^4 \times \frac{ 1 }{ x^2 }dx+c\] \[\frac{ y }{ x^2 }=\int\limits x^2 ~dx\] ?

OpenStudy (stamp):

I get \[y=x^6/4+Cx^2\] Solution claims\[2y=x^4+Cx^2\]

OpenStudy (stamp):

ohhhh hang on

OpenStudy (stamp):

ok i realied I wasnt mulitplying the right side by the integrating factor so now i get \[y=x^4/2+Cx^2\]

OpenStudy (stamp):

gottt it thx

OpenStudy (sshayer):

correction \[\frac{ y }{ x^2 }=\int\limits x^3\times \frac{ 1 }{ x^2 }dx=\frac{ x^2 }{ 2 }+c\] \[\frac{ 2 y }{ x^2 }=x^2+2c=x^2+ C\] \[2 y=x^4+Cx^2\] so you are correct.